# Personal LLM Agents: Insights and Survey about the Capability, Efficiency and Security

### v1

@@ Abstract
Since the advent of personal computing devices, intelligent personal assistants (IPAs) have been one of the key technologies that researchers and engineers have focused on, aiming to help users efficiently obtain information and execute tasks, and provide users with more intelligent, convenient, and rich interaction experiences. With the development of the smartphone and Internet of Things, computing and sensing devices have become ubiquitous, greatly expanding the functional boundaries of intelligent personal assistants. However, due to the lack of capabilities such as user intent understanding, task planning, tool using, and personal data management etc., existing intelligent personal assistants still have limited practicality and scalability.
In recent years, the emergence of foundation models, represented by large language models (LLMs), brings new opportunities for the development of intelligent personal assistants. With the powerful semantic understanding and reasoning capabilities, LLM can enable intelligent agents to solve complex problems autonomously. In this paper, we focus on Personal LLM Agents, which are LLM-based agents that are deeply integrated with personal data and personal devices and used for personal assistance. We envision that Personal LLM Agents will become a major software paradigm for end-users in the upcoming era. To realize this vision, we take the first step to discuss several important questions about Personal LLM Agents, including their architecture, capability, efficiency and security. We start by summarizing the key components and design choices in the architecture of Personal LLM Agents, followed by an in-depth analysis of the opinions collected from domain experts. Next, we discuss several key challenges to achieve intelligent, efficient and secure Personal LLM Agents, followed by a comprehensive survey of representative solutions to address these challenges.
@@ 2.2.2 Supervised Learning Methods
The idea of learning an interaction model from human interaction traces is introduced in Humanoid [30], which aims to generate human-like test inputs based on the GUI layout information. Seq2act [4] firstly focused on the mobile UI task automation domain, where the natural language instructions need to be mapped to a sequence of actions that can be directly executed. The framework decomposed the problem into an action phrase-extraction part and a grounding part, both using the Transformer [31] network. Inspired by the success of pretraining in NLP, ActionBert [32] uses self-supervised pretraining to enhance the model’s understanding of UIs. Specifically, to capture the semantics information of the UI switching actions, the model is designed to take a pair of UIs as input, and output embeddings of both UIs and individual components. Aimed at better compatibility with the restricted resource on mobile devices, Versatile UI Transformer (VUT) [33] was proposed to learn different UI grounding tasks within a single small model. It handles images, structures, and text-based types of data, using 3 task heads to support performing 5 distinct tasks simultaneously, including UI object detection, natural language command grounding, widget captioning, screen summarization and UI tappability prediction. Based on the self-aligned characteristics between components of different modalities, UIBert [34] presented a well-designed joint image-text model to utilize the correspondence, learning contextual UI embeddings from unlabeled data. To address the problem of lacking UI metadata, such as DOM tree and view hierarchy, SpotLight [35] introduced a vision-only approach for mobile UI understanding by taking screenshots and a region of interest (the “focus”) as input. Composed of a vision encoder and a language decoder, it can complete tasks according to the provided screenshot and prompt. Besides, Lexi [36] was proposed to leverage text-based instruction manuals and user guides to curate a multimodal dataset. By fusing text and visual features as input to the co-attention transformer layers, the model is pre-trained to form connections between text-based instructions and UI screenshots. UINav [37] utilized a referee model to evaluate the performance of the agent, immediately inform the users of the feedback. It also adopted demonstration augmentation to increase the data diversity.
@@ 2.2.3 Reinforcement Learning Methods
Another challenge of RL-based task automation is the huge action space and the sparse reward. A typical GUI-grounded task usually involves $5$-$10$ steps, each of which contains $10$-$100$ candidate actions, leading to a search space size of $10^{5}$-$100^{10}$. The task is completed only if the correct sequence of actions is taken. In order to tackle such challenge, many frameworks have been proposed. Liu et al. [6] introduced the method to use high-level “workflows” to constrain the allowable actions at each time step. The workflows can prune out bad exploration directions, accelerating the agent’s ability to discover rewards. Gur et al. [39] decomposed the complicated instruction into multiple smaller ones, and schedule a curriculum for the agents to gradually manage to follow an increasing number of sub-instructions. Besides, a meta-learning framework is also proposed to generate instruction-following tasks. Jia et al. [40] framed the actions of agent on the web into three distince categories, namely, DOM selection, token selection, and mode selection. What’s more, a factorized Q-value function is designed, assuming the independence of DOM selection and token selection. Glider [5] achieves its goal of reducing action space with a hierachical policy, which contains a master policy to handle the overall navigation and sub-policies to deal with specific widgets. Humphreys et al. [41] proposed the framework to directly use mouse and keyboard to complete tasks instead of depending on the specialized action spaces, which simplifies the use of behavioural priors informed by actual human-computer interactions
@@ 3.2 Intelligence Levels of Personal LLM Agents
L1 - Simple Step Following
Agent completes tasks by following exact steps predefined by the users or the developers.
- User: “Open Messenger”; Agent opens the app named Messenger.
@@
L2 - Deterministic Task Automation
Based on the user’s description of a deterministic task, agent auto-completes the necessary steps in a predefined action space.
- User: “Check the weather in Beijing today”; Agent automatically calls the weather API with parameter “Beijing” and parses info. from the response.
@@
L3 - Stratigic task Automation
Based on user-specified tasks, agents autonomously plan the execution steps using various resources and tools, and iterates the plan based on intermediate feedback until completion.
- User: “Tell Alice about my schedule for tomorrow”; Agent gathers tomorrow’s schedule information from the user’s calendar and chat history, then summarizes and sends them to Alice via Messenger.
@@
L4 - Memory and Context Awareness
Agent senses user context, understands user memory, and proactively provides personalized services at appropriate times.
- Agent recommends suitable financial products automatically based on User’s recent income and expenses, considering User’s personality and risk preference.
@@
L5 - Digital Persona
Agent fully represents the user in completing complex affairs, can interact on behalf of user with other users or agents, ensuring safety and reliability.
- Agent automatically reads emails and messages on behalf of User, replies to questions without user intervention, and summarizes them into an abstract.
@@
With ultra intelligence at level 5, agents play the role of a Digital Persona that can fully represent the user in completing complex affairs, thus users only need to focus on creativity and emotion. Agents not only sense the current status, but also predict the users’ future activities and take actions to facilitate them. Beyond directly serving users, Digital Persona can also collaborate with other agents to alleviate the burden of their users’ communication. Moreover, the level-5 agents should be able to continuously improve themselves through self-evolution.
@@ 3.3 Opinions on Common Problems
1
If the LLM is applied to personal intelligent agents, do you think it should be deployed locally or remotely?
2
How do you think customized models tailored for different users or organizations should be implemented?
3
For the LLM deployed on personal devices, which modality(ies) do you think needs to be supported?
4
What do you think is the most important capability of LLMs for personal LLM agents?
5
Considering the industry you are in, which ways of interaction do you think are the most promising for personal LLM agents?
6
In the future development of personal LLM agents, which aspect is the most crucial?
7
What features do you hope a future personal LLM agent can provide for you or your customers?
8
When integrating LLM with personal devices, what challenges do you think will be faced? What are the most urgent technical issues that needs to be addressed?
Opinion 1 (where to deploy the LLM): Edge-cloud (local-remote) collaborated deployment of LLM is preferred, while existing cloud-only (remote-only) (e.g.,, ChatGPT) is not a widely acceptable solution. As shown in Figure 7, 88% of participants prefer an edge-cloud collaborated architecture, 58.33% of them support local deployment, and 81.82% of them are not satisfied with the existing cloud-only solutions. Their main concerns are 1) the high latency of remote LLM service, 2) the privacy issue of transmitting personal data to the cloud, and 3) the huge cost of cloud-based LLM services.
@@ 4.1.1 Task Automation Methods
  • Program synthesis method is to utilize the code generation ability of LLMs to interact with APIs. One way is to fine-tune LLMs to use specific APIs. WebGPT [46] fine-tunes a GPT-3 [77] to answer long-form questions by calling Microsoft Bing Web Search API [78]. Some recent works [45, 79, 80, 81] fine-tune LLMs to retrieve and call APIs, enhancing their performance in various tasks like mathematical reasoning and program synthesis. Another way is to utilize the chain reasoning [82, 83, 67] and in-context learning ability [77] of LLMs. They show descriptions and demonstrations of the tools (e.g. APIs, other DNNs, etc.) in context and ask LLMs how to use them to complete tasks [84, 85, 86, 51, 87]. However, fine-tuning LLMs can be costly and restricted to the predefined set of tools, and in-context learning may fail when the number of APIs go large. Thus, authors of ToolkenGPT [88] attempt to solve this problem by representing each tool (API) as a token.
@@
  • Text-based GUI representation is to convert the UIs into pure text. Seq2act [4] uses a transformer-based model [31] to ground users’ instruction to UI actions described in <operation, object, argument> tuples. Researchers also investigate prompting with mobile UIs to complete tasks of UI instruction mapping [89]. The authors convert mobile UI into HTML code, which is easy for LLMs to understand because an important part of their training data is scraped from Github. DroidBot-GPT [90] is an LLM-based system to complete users’ tasks in a sequence of UI actions. Mind2Web [91] filters the raw HTML of webpages with a smaller LM and uses the LLM to select the target element and action. AutoDroid [92] uses app analysis tools to acquire app domain-specific knowledge and uses it to augment the LLMs for task automation. In AXNav [93], authors build a system using LLMs and pixel-based UI Understanding to execute manual accessibility tests. MemoDroid [94] introduces an LLM-based mobile task automator that can break tasks into smaller sub-tasks and complete them by recalling former actions.
  • Multimodal representation is to use the image (and text) description of UI as the input of the Personal LLM Agents. META-GUI [95] proposes a UI-based method for task-oriented dialogue (TOD) systems. ResponsibleTA [96] proposes a fundamental multi-modal framework that equips LLMs with the ability to predict feasibility, verify completeness, and ensure security when autonomously completing users’ tasks. Auto-UI [97] proposes a Multi-modal framework that interacts directly with the interface and introduces a chain-of-action technique to aid the agent in making decisions. RUIG [98] builds a multimodal model to map instructions to UI screenshots, and MM-Navigator [99] uses GPT4-V for UI navigation tasks. UINav [37] introduces a lightweight task automation system based on task demonstrations, and the input modality can dynamically switch between raw pixels view hierarchy. CogAgent [100] introduces a visual language model with 18 billion parameters to complete tasks such as UI understanding and navigation.
@@ 4.1.3 Evaluation
Benchmarks: Table 6 lists the benchmarks of UI-based task automation. One group of benchmarks is static datasets, which often include a set of human-annotated tasks, structured UI data (and screenshots), and actions to complete the tasks. Some of the tasks are synthetically generated [4, 116, 117]. The early works mainly focus on low-level tasks with clear instructions [118, 34], for example, click the ‘settings’ button, and then click ‘Font size’. Later works introduce high-level tasks that could be completed in multiple steps [4, 119, 95, 120, 91, 121], for example, delete all the events in my calendar. Another group of benchmarks are platforms that enable the agent to interact with. MiniWoB++ [38, 6], WebShop [122], and WebArena [123] provide web environments where agents can navigate and operate on the web by clicking, typing, closing page, and so on. AndroidEnv [116] and MobileEnv [117] provide a dynamic environment where agents can engage with any Android-based application and the core operating system. This framework allows for a wide scope of interaction and task-solving capabilities within the diverse Android platform.
@@ 4.2.1 Sensing Sources
Software Sensor. Unlike hardware sensing that obtains data from real sensor devices, software sensing focuses on obtaining information from existing data, such as app usage [131], call records [132], and typing habits [133], etc. In reality, the realm of software sensing is incredibly broad. For instance, in the field of natural language processing or audio, there exists a plethora of sensing research based on text or speech. Furthermore, recommendation systems such as e-commerce or short video platforms, the process typically involves first sensing certain user information and subsequently recommending specific products or content. These sensors let agents better understand the users, enabling them to provide with more intelligent and personalized services.

### v2

@@ Abstract
Since the advent of personal computing devices, intelligent personal assistants (IPAs) have been one of the key technologies that researchers and engineers have focused on, aiming to help users efficiently obtain information and execute tasks, and provide users with more intelligent, convenient, and rich interaction experiences. With the development of the smartphone and Internet of Things, computing and sensing devices have become ubiquitous, greatly expanding the functional boundaries of IPAs. However, due to the lack of capabilities such as user intent understanding, task planning, tool using, and personal data management etc., existing IPAs still have limited practicality and scalability.
Recently, the emergence of foundation models, represented by large language models (LLMs), brings new opportunities for the development of IPAs. With the powerful semantic understanding and reasoning capabilities, LLM can enable intelligent agents to solve complex problems autonomously. In this paper, we focus on Personal LLM Agents, which are LLM-based agents that are deeply integrated with personal data and personal devices and used for personal assistance. We envision that Personal LLM Agents will become a major software paradigm for end-users in the upcoming era. To realize this vision, we take the first step to discuss several important questions about Personal LLM Agents, including their architecture, capability, efficiency and security. We start by summarizing the key components and design choices in the architecture of Personal LLM Agents, followed by an in-depth analysis of the opinions collected from domain experts. Next, we discuss several key challenges to achieve intelligent, efficient and secure Personal LLM Agents, followed by a comprehensive survey of representative solutions to address these challenges.
@@ 2.2.2 Supervised Learning Methods
The idea of learning an interaction model from human interaction traces is introduced in Humanoid [30], which aims to generate human-like test inputs based on the GUI layout information. Seq2act [4] firstly focused on the mobile UI task automation domain, where the natural language instructions need to be mapped to a sequence of actions that can be directly executed. The framework decomposed the problem into an action phrase-extraction part and a grounding part, both using the Transformer [31] network. Inspired by the success of pretraining in NLP, ActionBert [32] uses self-supervised pretraining to enhance the model’s understanding of UIs. Specifically, to capture the semantics information of the UI switching actions, the model is designed to take a pair of UIs as input, and output embeddings of both UIs and individual components. Fu et al. [33] extended the concept of Words/Sentences from NLP to Pixel-Words/Screen-Sentences. By pre-training with visual atomic components (Pixel-Words), the PW2SS framework (Sentence Transformer) could accomplish various downstream GUI understanding tasks. Aimed at better compatibility with the restricted resource on mobile devices, Versatile UI Transformer (VUT) [34] was proposed to learn different UI grounding tasks within a single small model. It handles images, structures, and text-based types of data, using 3 task heads to support performing 5 distinct tasks simultaneously, including UI object detection, natural language command grounding, widget captioning, screen summarization and UI tappability prediction. Based on the self-aligned characteristics between components of different modalities, UIBert [35] presented a well-designed joint image-text model to utilize the correspondence, learning contextual UI embeddings from unlabeled data. To address the problem of lacking UI metadata, such as DOM tree and view hierarchy, SpotLight [36] introduced a vision-only approach for mobile UI understanding by taking screenshots and a region of interest (the “focus”) as input. Composed of a vision encoder and a language decoder, it can complete tasks according to the provided screenshot and prompt. Besides, Lexi [37] was proposed to leverage text-based instruction manuals and user guides to curate a multimodal dataset. By fusing text and visual features as input to the co-attention transformer layers, the model is pre-trained to form connections between text-based instructions and UI screenshots. UINav [38] utilized a referee model to evaluate the performance of the agent, immediately inform the users of the feedback. It also adopted demonstration augmentation to increase the data diversity.
@@ 2.2.3 Reinforcement Learning Methods
Another challenge of RL-based task automation is the huge action space and the sparse reward. A typical GUI-grounded task usually involves $5$-$10$ steps, each of which contains $10$-$100$ candidate actions, leading to a search space size of $10^{5}$-$100^{10}$. The task is completed only if the correct sequence of actions is taken. In order to tackle such challenge, many frameworks have been proposed. Liu et al. [6] introduced the method to use high-level “workflows” to constrain the allowable actions at each time step. The workflows can prune out bad exploration directions, accelerating the agent’s ability to discover rewards. Gur et al. [40] decomposed the complicated instruction into multiple smaller ones, and schedule a curriculum for the agents to gradually manage to follow an increasing number of sub-instructions. Besides, a meta-learning framework is also proposed to generate instruction-following tasks. Jia et al. [41] framed the actions of agent on the web into three distince categories, namely, DOM selection, token selection, and mode selection. What’s more, a factorized Q-value function is designed, assuming the independence of DOM selection and token selection. Glider [5] achieves its goal of reducing action space with a hierachical policy, which contains a master policy to handle the overall navigation and sub-policies to deal with specific widgets. Humphreys et al. [42] proposed the framework to directly use mouse and keyboard to complete tasks instead of depending on the specialized action spaces, which simplifies the use of behavioural priors informed by actual human-computer interactions.
@@ 3.2 Intelligence Levels of Personal LLM Agents
L1 - Simple Step Following Agent completes tasks by following exact steps predefined by the users or the developers. - User: “Open Messenger”; Agent opens the app named Messenger.
@@
L2 - Deterministic Task Automation Based on the user’s description of a deterministic task, agent auto-completes the necessary steps in a predefined action space. - User: “Check the weather in Beijing today”; Agent automatically calls the weather API with parameter “Beijing” and parses info. from the response.
@@
L3 - Strategic task Automation Based on user-specified tasks, agents autonomously plan the execution steps using various resources and tools, and iterates the plan based on intermediate feedback until completion. - User: “Tell Alice about my schedule for tomorrow”; Agent gathers tomorrow’s schedule information from the user’s calendar and chat history, then summarizes and sends them to Alice via Messenger.
@@
L4 - Memory and Context Awareness Agent senses user context, understands user memory, and proactively provides personalized services at appropriate times. - Agent recommends suitable financial products automatically based on User’s recent income and expenses, considering User’s personality and risk preference.
@@
L5 - Autonomous Avatar Agent fully represents the user in completing complex affairs, can interact on behalf of user with other users or agents, ensuring safety and reliability. - Agent automatically reads emails and messages on behalf of User, replies to questions without user intervention, and summarizes them into an abstract.
@@
With ultra intelligence at level 5, agents play the role of an Autonomous Avatar that can fully represent the user in completing complex affairs, thus users only need to focus on creativity and emotion. Agents not only sense the current status, but also predict the users’ future activities and take actions to facilitate them. Beyond directly serving users, an Autonomous Avatar can also collaborate with other agents to alleviate the burden of their users’ communication. Moreover, the level-5 agents should be able to continuously improve themselves through self-evolution.
@@ 3.3 Opinions on Common Problems
1 If the LLM is applied to personal intelligent agents, do you think it should be deployed locally or remotely?
2 How do you think customized models tailored for different users or organizations should be implemented?
3 For the LLM deployed on personal devices, which modality(ies) do you think needs to be supported?
4 What do you think is the most important capability of LLMs for personal LLM agents?
5 Considering the industry you are in, which ways of interaction do you think are the most promising for personal LLM agents?
6 In the future development of personal LLM agents, which aspect is the most crucial?
7 What features do you hope a future personal LLM agent can provide for you or your customers?
8 When integrating LLM with personal devices, what challenges do you think will be faced? What are the most urgent technical issues that needs to be addressed?
Opinion 1 (where to deploy the LLM): Edge-cloud (local-remote) collaborated deployment of LLM is preferred, while existing cloud-only (remote-only) (e.g., ChatGPT) is not a widely acceptable solution. As shown in Figure 7, 88% of participants prefer an edge-cloud collaborated architecture, 58.33% of them support local deployment, and 81.82% of them are not satisfied with the existing cloud-only solutions. Their main concerns are 1) the high latency of remote LLM service, 2) the privacy issue of transmitting personal data to the cloud, and 3) the huge cost of cloud-based LLM services.
@@ 4.1.1 Task Automation Methods
  • Program synthesis method is to utilize the code generation ability of LLMs to interact with APIs. One way is to fine-tune LLMs to use specific APIs. WebGPT [47] fine-tunes a GPT-3 [78] to answer long-form questions by calling Microsoft Bing Web Search API [79]. Some recent works [46, 80, 81, 82] fine-tune LLMs to retrieve and call APIs, enhancing their performance in various tasks like mathematical reasoning and program synthesis. Octopus V2 [83] introduces a 2B parameter on-device LLM to call Android APIs for task automation. Another way is to utilize the chain reasoning [84, 85, 68] and in-context learning ability [78] of LLMs. They show descriptions and demonstrations of the tools (e.g. APIs, other DNNs, etc.) in context and ask LLMs how to use them to complete tasks [86, 87, 88, 52, 89]. However, fine-tuning LLMs can be costly and restricted to the predefined set of tools, and in-context learning may fail when the number of APIs go large. Thus, authors of ToolkenGPT [90] attempt to solve this problem by representing each tool (API) as a token.
@@
  • Text-based GUI representation is to convert the UIs into pure text. Seq2act [4] trains a transformer-based model [31] to ground users’ instruction to UI actions described in <operation, object, argument> tuples. Researchers also investigate prompting with mobile UIs to complete tasks of UI instruction mapping [91]. The authors convert mobile UI into HTML code, which is easy for LLMs to understand because an important part of their training data is scraped from Github. DroidBot-GPT [92] is an LLM-based system to complete users’ tasks in a sequence of UI actions. Mind2Web [93] filters the raw HTML of webpages with a smaller LM and uses the LLM to select the target element and action. AutoDroid [94] uses app analysis tools to acquire app domain-specific knowledge and uses it to augment the LLMs for task automation. In AXNav [95], authors build a system using LLMs and pixel-based UI Understanding to execute manual accessibility tests. MemoDroid [96] introduces an LLM-based mobile task automator that can break tasks into smaller sub-tasks and complete them by recalling former actions.
  • Multimodal representation is to use the image (and text) description of UI as the input of the Personal LLM Agents. Early research work is focused on training multimodal transformers to ground user commands to UI elements [97, 98, 38]. In the era of LLMs, some approaches attempted to combine visual encoders with LLMs to handle GUI images [99, 100, 101]. With the advent of Large Multimodal Models (LMMs), a growing number of projects employed visual language agents for UI action grounding and navigation [102, 103]. One trend involves leveraging powerful LMMs such as GPT-4V to comprehend GUIs and select UI elements [104, 105, 106, 107]. Another line of research is to customize open-sourced LMMs by fine-tuning on large-scale datasets for GUI-related tasks [108, 109, 110].
@@ 4.1.3 Evaluation
OmniACT [132] Desktop+Web 9,802 VH, Screen
AutoWebBench [133] Web 10,000 DOM, Screen
VisualWebBench [134] Web 1,500 DOM, Screen
ScreenAgent [135] Desktop 273 Screen
@@
OSWorld [103] Desktop 369 VH, Screen
AgentStudio [138] Desktop+Web 227 DOM, Screen
Benchmarks: Table 6 lists the benchmarks of UI-based task automation. One group of benchmarks is static datasets, which often include a set of human-annotated tasks, structured UI data (and screenshots), and actions to complete the tasks. Some of the tasks are synthetically generated [4, 126, 127]. The early works mainly focus on low-level tasks with clear instructions [128, 35], for example, click the ‘settings’ button, and then click ‘Font size’. Later works introduce high-level tasks that could be completed in multiple steps [4, 129, 97, 130, 93, 131, 132, 133, 134, 135], for example, delete all the events in my calendar. Another group of benchmarks are platforms that enable the agent to interact with. MiniWoB++ [39, 6], WebShop [136], and WebArena [137] provide web environments where agents can navigate and operate on the web by clicking, typing, closing page, and so on. AgentStudio [138] provides a comprehensive platform that supports interactions with versatile real-world computers. AndroidEnv [126] and MobileEnv [127] provide a dynamic environment where agents can engage with any Android-based application and the core operating system. This framework allows for a wide scope of interaction and task-solving capabilities within the diverse Android platform.
@@ 4.2.1 Sensing Sources
Recently, there has been a proliferation of research exploring the deep integration of LLMs with raw sensor data. For instance, several studies directly embed raw IMU data into prompts for LLM, enabling Human Activity Recognition (HAR) [145] or trajectory prediction [146]. Zhang et al. [147] provides LLM with a bird’s-eye view of a 3D scene and allows it to iteratively select viewpoints to understand 3D point cloud scenes. Additionally, Zheng et al. [148] employs a trainable dual-channel audio frontend and fine-tuned LLM to enable LLM to comprehend spatial sound. Similar frontend and fine-tuning approaches are prevalent in various domains such as LiDAR [149] and autonomous driving [150, 151].
Software Sensor. Unlike hardware sensing that obtains data from real sensor devices, software sensing focuses on obtaining information from existing data, such as app usage [152], call records [153], typing habits [154], video game [155], etc. The scope of software sensing is incredibly broad. For instance, in the field of natural language processing or audio, there exists a plethora of sensing research based on text or speech. Furthermore, recommendation systems such as e-commerce or short video platforms, the process typically involves first sensing certain user information and subsequently recommending specific products or content. These sensors let agents better understand the users, enabling them to provide with more intelligent and personalized services.
@@
There are three different approaches to enable LLM to understand and utilize sensor data.
  • Option 1: Sensor Data as Prompt. This method directly inputs sensor data into LLM as text prompts. Such an approach can be applied to various sensing sources such as IMU [146] and bluetooth [164]. The mappings between the raw sensor data and the prompts can be created through rules, such as mapping tactile sensations on object surfaces to descriptors like “soft” or “hard” [165]. This method is simple and effective as demonstrated many existing studies. However, it also has important limitations, such as the significant computational cost of processing large volumes of raw data and the limited ability of LLM to understand the complex sensor data in plain text.
  • Option 2: Sensor Data Encoding + Fine-tuning. This approach enables LLM to understand sensor data with a data encoder. The encoder generates token embeddings from the raw sensor data with a learned neural network, and the embeddings are usually integrated into LLM through fine-tuning. This method yields significant results for complex sensor data, such as LiDAR [149] and dual-channel audio [148]. This approach allows LLM to efficiently understand sensor modalities, which is used to construct complex end-to-end systems like autonomous driving [151, 150]. Its drawback lies in the high training difficulty.
  • Option 3: Redirecting Sensor Data to Domain-Specific Models. This approach doesn’t process sensor data directly with LLMs, while it uses LLMs to invoke other specialized small models to deal with the raw sensor data. For example, Darvish et al. [166] leveraging techniques like object detection or pose estimation to assist chemical experiment robots in improving perception and understanding, additional information is added to the raw data stream and transformed into a form that LLMs can understand.